

********************************************************************
UARS WEEKLY STATUS REPORT - 13 MAY 2005
********************************************************************
***********
SOLSTICE
***********
This week SOLSTICE has been under daylight only operations due to low
power availability. Otherwise, the instrument is in good health
gathering solar data on all available orbital tracking periods. On
average, we received 73% of each day's data amounting to about 388
minutes of solar observing time per day.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005029 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2004366 has been processed and is waiting for DAAC archival.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 30 April 2005 (Orbit 74625, GMT Day 120) through
06 May 2005 (Orbit 74729, GMT Day 126).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 126/1401 74724 POWER ON
The beta angle ranged between the angles of 04.3 degrees to 25.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 38 degrees on 14 May 2005 (DOY 134).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 7.20 to 7.09 degrees C. Additional
battery performance data for each battery on 29 April 2005 and 06 May
2005 are provided in Appendix D.
The clock error ranged from 10.7 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 06
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the status of the
batteries. The FOT met with the NASA power branch engineers and
decided that the HALOE turn on would be delayed by two days to attempt
to fully charge battery 3. This was done and the HALOE turn-on was
successfully executed on 06 May 2005. Subsequent to the HALOE turn-
on, it was determined that VT-6 will be maintained until the battery 2
temperature or C/D ratio increases above their nominal range.
Also this week, the FOT continued its testing of the solar array
drive, with the goal of finding a way to again neutralize the A-side.
So far this has not been accomplished, but testing will continue next
week.
MEETINGS
The FOT met with the NASA power branch engineers on 02 May 2005 to
discuss the battery configuration plan for the upcoming week.
The FOT also attended the UARS Monthly Status review on 03 May 2005 to
review the significant events from the month of April.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 30 May 2005 (DOY 150). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 May 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
74630 120/0805 05-158 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
74666 122/1812 05-159 SADDE TESTING THRU ORBIT 74667
74687 124/0427 05-160 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
74693 124/1305 05-161 SADDE TESTING THRU ORBIT 74697
74697 124/1956 05-162 NBTR MUB INHIBIT
74706 125/0939 05-163 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
74724 126/1401 05-164 HALOE POWER ON - PART ONE
74726 126/1834 05-165 HALOE POWER ON - PART TWO
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 May 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 May 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
74630 120/0805 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
74687 124/0433 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
74706 125/0939 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 May 2005
29 April 2005 - GMT Day 119 Beta = 01.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -50.4/-078.4 -3.68/-4.94 +00.0/+00.0 06.1 0.00 0.0
2 -67.2/-112.0 +7.74/+6.11 +33.6/-18.0 26.4 1.03 22.72
3 235.2/-700.0 +2.60/+1.34 +18.8/-10.4 26.2 1.06 11.07
06 May 2005 - GMT Day 126 Beta = 25.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -22.4/ -78.4 -3.05/-4.31 +00.0/+00.0 05.8 0.00 0.0
2 179.2/ -05.6 +11.07/+9.06 +35.6/-20.4 25.8 1.05 24.62
3 207.2/-700.0 +3.87/+1.97 +17.6/-11.2 25.8 1.06 12.16
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 May 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-023 FOT 74635 120/1601 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-026 SOLSTICE 74520 113/0053 N/A SOLS TEMP LIMITS
05-027 FOT 74548 114/2224 39337 Data Drop Out
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-023 FOT See List Apr 05 30832 Generic Late Acquisition